# Limits of Probability Measures with General Coefficients

### v2

@@ 2.2. Constant term
The most important part of the formal series after applying the $\mathcal{Q}_{i}^{N}$ operators is the term with degree 00, see Section 3. For a sequence of coefficients $\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$, this corresponds to $c^{0,0}$. In this subsection, we show Theorem 2.9 to compute the degree 00 term following $\mathcal{Q}_{i}^{N}$ operators when there is only one value of $i$. Later on, in Theorem 2.13, we look at the constant term following $\mathcal{Q}_{i}^{N}$ operators when there are any number of distinct values of $i$. These values are computed with free cumulants, which are introduce in Definition 2.8. However, we must first introduce noncrossing partitions.
@@
with $B_{i}$, $1\leq i\leq m$ being the blocks of $\pi$, such that the smallest element of $B_{i+1}$ is greater than the smallest element of $B_{i}$ for $11\leq i\leq m-$ 1. Also, the length of $\pi$ is $\ell(\pi)=m$.
@@
Suppose that for $k\geq 1$, $NC^{\prime}(k)=NC(\{1,3,4,\ldots,k+1\})$. For all $\pi\in NC^{\prime}(k)$, define $S(\pi)$ to be the set of partitions $\lambda$ in $NC(k+1)$ such that $\lambda$ is $\pi$ with $2$ added to $B_{1}$ or $\lambda$ is $\pi$ with $B_{1}$ replaced by $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$, where $B_{1}^{\prime}$ is the $j-1$ largest elements of $B_{1}$ with $1$ added, $B_{1}^{\prime\prime}$ is the remaining elements of $B_{1}$ with $2$ added, and $1\leq j\leq|B_{1}|$.
@@
First, we must prove that if $\pi\in NC^{\prime}(k)$, then $S(\pi)\subset NC(k+1)$. It is clear that $\lambda\in S(\pi)$ that is $\pi$ with $2$ added to $B_{1}$ is in $NC(k+1)$. Suppose $\lambda\in S(\pi)$ is $\pi$ with $B_{1}$ replaced by $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$, where $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$ are defined as mentioned previously. The noncrossing condition for $\lambda$ is clearly satisfied when the two distinct blocks are $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$. When the two distinct blocks are both not $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$, the noncrossing condition is also satisfied since $\pi$ is noncrossing. When one block is $B_{1}^{\prime}$ and the other is not $B_{1}^{\prime\prime}$, the noncrossing condition is satisfied because $B_{1}^{\prime}\subset B_{1}$ and $\pi$ is noncrossing. On the other hand, when one block is $B_{1}^{\prime\prime}$ and the other is not $B_{1}^{\prime}$, the noncrossing condition is satisfied because $B_{1}^{\prime\prime}$ is a subset of $B_{1}\cup\{2\}\backslash\{1\}$ and $\pi$ is noncrossing.
@@
For $\lambda\in NC(k+1)$, suppose the decomposition into blocks is $\lambda=C_{1}\sqcup C_{2}\sqcup\cdots\sqcup C_{m}$, such that the smallest element of $C_{i+1}$ is greater than the smallest element of $C_{i}$ for $11\leq i\leq m-$ 1. Note that $2\in C_{1}$ or $2\in C_{2}$.
If $2\in C_{1}$, let $\pi=B_{1}\sqcup\cdots\sqcup B_{m}\in NC^{\prime}(k)$ have $B_{i}=C_{i}$ for $2\leq i\leq m$ and $B_{1}=C_{1}\backslash\{2\}$. We see that $\pi$ is noncrossing and adding $2$ to $B_{1}$ gives $\lambda$, so $\lambda\in S(\pi)$. Suppose $\lambda\in S(\pi)$, $\pi=B_{1}\sqcup\cdots\sqcup B_{m^{\prime}}\in NC^{\prime}(k)$. Since the first block of $\lambda$ contains a $2$ and the only element of $S(\pi)$ with first block containing $2$ is $\pi$ with $2$ added to $B_{1}$, we know that $\pi$ must be $\lambda$ with $2$ removed from $C_{1}$. So, $\pi$ is unique, and $\lambda$ is not in $S(\pi)$ for more than one $\pi\in NC^{\prime}(k)$.
@@
Suppose that $2\in C_{2}$. Then, let $\pi=B_{1}\sqcup\cdots\sqcup B_{m-1}\in NC^{\prime}(k)$ have $B_{1}=\left(C_{1}\cup C_{2}\right)\backslash\{2\}$ and $B_{i-1}=C_{i}$ for $3\leq i\leq m$. Since $\lambda$ is noncrossing, $\pi$ is also noncrossing. The argument for this is as follows. It is clear that the noncrossing condition between $B_{i}$ and $B_{j}$ is satisfied when $12\leq i<j\leq m-$ 1. It suffices to prove that the noncrossing condition between $B_{1}=(C_{1}\cup C_{2})\backslash\{2\}$ and $B_{i}=C_{i+1}$ is satisfied for $2\leq i\leq m$. If there exists $b\in B_{1}$ and $c,d\in B_{i}$ such that $c<b<d$, then $1<c<b<d$ and $2<c<b<d$, which contradictions the noncrossing condition between either $C_{1}$ and $C_{i+1}$ or $C_{2}$ and $C_{i+1}$. Hence, $\pi$ is noncrossing.
@@
Note that $1\in C_{1}$ and $2\in C_{2}$. Suppose $j_{1}\in C_{1}$ and $j_{2}\in C_{2}$, where $j_{1},j_{2}>2$. Because $\lambda$ is noncrossing, we must have that $j_{1}>j_{2}.$ Therefore, the elements of $C_{1}$ which are not $1$ are greater than all of the elements of $C_{2}$. We see that, where $1\leq j=|C_{1}|\leq|B_{1}|$, $C_{1}$ is the largest $j-1$ elements of $B_{1}$ with $1$ added, and $C_{2}$ is the remaining elements of $B_{1}$ with $2$ added. Therefore, $\lambda\in S(\pi)$.
@@
Assume $\lambda\in S(\pi)$, where $\pi=B_{1}\sqcup\cdots\sqcup B_{m^{\prime}}\in NC^{\prime}(k)$. Since $\lambda$ does not have $2$ in the first block, $\lambda$ is $\pi$ with $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$ instead of $B_{1}$, where $B_{1}^{\prime}$ is the last $j-1$ elements of $B_{1}$ with $1$ added, and $B_{1}^{\prime\prime}$ is the remaining elements of $B_{1}$ with $2$ added, for some $1\leq j\leq|B_{1}|$. Then, the first and second block of $\lambda$ would be $C_{1}=B_{1}^{\prime}$ and $C_{2}=B_{1}^{\prime\prime}$, respectively. But, in $\pi$, $B_{1}=(B_{1}^{\prime}\cup B_{1}^{\prime\prime})/\{2\}=(C_{1}\cup C_{2})/\{2\}$ and $B_{i}=C_{i+1}$ for $12\leq i\leq m-$ 1, where $m^{\prime}=m-1$. Thus, $\pi$ is unique. ∎
With 2.7.1,
@@
giving Lemma 2.7 since |$|\{1,3,4,\ldots,k+1\}|=|\{1,2,3,\ldots,k\}|=k$. We are done. ∎
@@
The free cumulants appear in (8) and these free cumulants will be used in Section 3 and Section 4 to obtain the free cumulants corresponding to the moments $m_{k}$ in Theorem 1.7 that we mentioned previously. Furthermore, Theorem 1.9 is based on the asymptotics of the free cumulants.
@@
Let $i$ and $k$ be positive integers. Suppose $\{f_{j,N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ for $11\leq j\leq k-$ 1 and $\{g_{N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ are sequences of formal series which are symmetric outside of $i$. For $11\leq j\leq k-$ 1, assume that $f_{j,N}(x_{1},\ldots,x_{N})$ has limiting sequence $f$ outside of $i$. Also, suppose $g_{N}(x_{1},\ldots,x_{N})$ has limiting sequence $g$ outside of $i$. Then,
@@
From Corollary 2.4, the limit of $\prod_{j=1}^{k-1}\mathcal{Q}_{i}^{N}(f_{j,N})g_{N}(x_{1},\ldots,x_{N})$ as $N\rightarrow\infty$ outside of $i$ exists, see the definition of being symmetric outside of $i$. Thus, the left hand side of (8) is
$[1]\lim_{N\rightarrow\infty}\prod_{j=1}^{k-1}\mathcal{Q}_{i}^{N}(f_{j,N})g_{N}(x_{1},\ldots,x_{N}),$
@@
Next, assume that the statement is true for a positve integer $k$. We want to show the statement is true for $k+1$. From Proposition 2.5, outside of $i$,
@@
Using Corollary 2.4, let $g^{\prime}$ be the limiting sequence of $\mathcal{Q}_{i}^{N}(F_{i}(f))F_{i}(g)(x_{1},\ldots,x_{N})$ outside of $i$, with
@@
Also, from Lemma 2.7,
@@
Fix the value of $a$. We use induction on $m$, from $m=a$ to 00. The $m=a$ base case is clear. Suppose the statement holds for $m+1$, $0\leq m<a$. We show the statement holds for $m$ with induction on $b$, where the base case $b=0$ is clear. For the inductive step, assume the statement for $m$, $a$, and $b\geq 0$ holds. We want to show the statement for $m$, $a$, and $b+1$ holds. Note that
@@
Afterwards, the coefficient of $c_{f}^{d_{1},\nu_{1}}c_{g}^{d_{2},\nu_{2}}$ on the left hand side of (10) is, with 2.9.1 and 2.9.2,
@@ 2.3. Distinct Indices
Also, we look at the constant term after using combinations of $\mathcal{Q}_{i}^{j}$ with distinct indices $i$ in Theorem 2.13. This is important for the proof of Theorem 1.7, as will be seen in Section 3 and Section 4, and for this, we use the operator in (11). But, it is necessary to consider the free cumulants of $s=\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$ which are symmetric in the $x_{i}$ beforehand.
@@
By Corollary 2.4, the resulting formal series has variables $x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{j}$ and is symmetric.
@@
Induction on $m$ is used. For the base case $m=1$, $\lim_{N\rightarrow\infty}\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ exists outside of $1$ from Corollary 2.4 because $s_{1,N}f_{1,N}$ and $s_{1,N}g_{N}$ have limits outside of $1$. For some $i\geq 1$, consider $N\geq i+1$. From Lemma 2.11, $\mathcal{R}_{N,N}^{\lambda_{1}}(f_{1,N})(g_{N})\in\mathcal{F}_{i}^{N-1}$.
From Proposition 2.6, $s_{1,N}\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})=\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$. With this, the coefficient $c^{d,\nu}$ of $\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})|_{x_{N}=0}$ as a formal series in $x_{1},\ldots,x_{N-1}$ outside of $i$ will be the coefficient $c^{0,\nu+(d)}$ of $\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})$ outside of $N$, and thus the coefficient $c^{0,\nu+(d)}$ of $\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ outside of $1$. Since the limit of $\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ outside of $1$ exists, the limit of this coefficient from $N=i+1$ to $\infty$ exists. As this holds for all $d,\nu$, the base case is complete.
@@
By the inductive hypothesis and because $h_{N-m}$ is symmetric by Lemma 2.11, the limit of $s_{1,N-m}h_{N-m}=h_{N-m}$ from $N-m=1$ to $\infty$ exists outside of $1$. Then,
@@
is the base case $m=1$ with $f_{m+1,N}$ as $f_{1,N-m}$, $h_{N-m}$ as $g_{N-m}$, and $N-m$ as $N$, where the conditions are satisfied. Therefore, from $m=1$, the limit from $N-m=i+1$ to $\infty$ of the above expression outside of $i$ exists for $i\geq 1$. This limit is from $N=i+m+1$ to $\infty$, which completes the proof. ∎
@@
We prove this with induction on $m$. For the base case $m=1$, use Theorem 2.9. Suppose the result holds for $m\geq 1$. We want to show it holds for $m+1$.
@@
Then, $s_{1,N-m}\mathcal{Q}_{N-m}^{N-m}\left(f_{m+1,N}\right)^{\lambda_{m+1}}h_{N-m}$ is $\mathcal{Q}_{1}^{N-m}\left(s_{1,N-m}f_{m+1,N}\right)^{\lambda_{m+1}}h_{N-m}$ from Proposition 2.6. Since the switch does not change the constant,
@@
The limit of this as $N\rightarrow\infty$ is computed. With $f_{N-m}=s_{1,N-m}f_{m+1,N}$ for $N\geq m+1$, $f_{N}$ has $N\rightarrow\infty$ limit $f_{m+1}$ outside of $1$, and from Lemma 2.12, $h_{N}$ has a limit outside of $1$; let $h$ be the limiting sequence of $\{h_{N}\}_{N\geq 1}$. Then, from Theorem 2.9 with $N-m$ for $N$,
@@
But, as $h_{N}$ is symmetric in $x_{i}$, $1\leq i\leq N$ for $N\geq 1$, $h$ is symmetric. Then, by Proposition 2.10, $c_{k}(h)=0$, $k\geq 2$, and $c_{1}(h)=c_{h}^{0,0}$. From this,

### v3

@@ 2.2. Constant term
The most important part of the formal series after applying the $\mathcal{Q}_{i}^{N}$ operators is the term with degree 0, see Section˜3. For a sequence of coefficients $\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$, this corresponds to $c^{0,0}$. In this subsection, we show Theorem˜2.9 to compute the degree 0 term following $\mathcal{Q}_{i}^{N}$ operators when there is only one value of $i$. Later on, in Theorem˜2.13, we look at the constant term following $\mathcal{Q}_{i}^{N}$ operators when there are any number of distinct values of $i$. These values are computed with free cumulants, which are introduce in Definition˜2.8. However, we must first introduce noncrossing partitions.
@@
with $B_{i}$, $1\leq i\leq m$ being the blocks of $\pi$, such that the smallest element of $B_{i+1}$ is greater than the smallest element of $B_{i}$ for $1\leq i\leq m-1$. Also, the length of $\pi$ is $\ell(\pi)=m$.
@@
Suppose that for $k\geq 1$, $NC^{\prime}(k)=NC(\{1,3,4,\ldots,k+1\})$. For all $\pi\in NC^{\prime}(k)$, define $S(\pi)$ to be the set of partitions $\lambda$ in $NC(k+1)$ such that $\lambda$ is $\pi$ with $2$ added to $B_{1}$ or $\lambda$ is $\pi$ with $B_{1}$ replaced by $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$, where $B_{1}^{\prime}$ is the j1j-1 largest elements of $B_{1}$ with $1$ added, $B_{1}^{\prime\prime}$ is the remaining elements of $B_{1}$ with $2$ added, and $1\leq j\leq|B_{1}|$.
@@
[…]
@@
For $\lambda\in NC(k+1)$, suppose the decomposition into blocks is $\lambda=C_{1}\sqcup C_{2}\sqcup\cdots\sqcup C_{m}$, such that the smallest element of $C_{i+1}$ is greater than the smallest element of $C_{i}$ for $1\leq i\leq m-1$. Note that $2\in C_{1}$ or $2\in C_{2}$.
[…]
@@
Suppose that $2\in C_{2}$. Then, let $\pi=B_{1}\sqcup\cdots\sqcup B_{m-1}\in NC^{\prime}(k)$ have $B_{1}=\left(C_{1}\cup C_{2}\right)\backslash\{2\}$ and $B_{i-1}=C_{i}$ for $3\leq i\leq m$. Since $\lambda$ is noncrossing, $\pi$ is also noncrossing. The argument for this is as follows. It is clear that the noncrossing condition between $B_{i}$ and $B_{j}$ is satisfied when $2\leq i<j\leq m-1$. It suffices to prove that the noncrossing condition between $B_{1}=(C_{1}\cup C_{2})\backslash\{2\}$ and $B_{i}=C_{i+1}$ is satisfied for $2\leq i\leq m$. If there exists $b\in B_{1}$ and $c,d\in B_{i}$ such that $c<b<d$, then $1<c<b<d$ and $2<c<b<d$, which contradictions the noncrossing condition between either $C_{1}$ and $C_{i+1}$ or $C_{2}$ and $C_{i+1}$. Hence, $\pi$ is noncrossing.
@@
Note that $1\in C_{1}$ and $2\in C_{2}$. Suppose $j_{1}\in C_{1}$ and $j_{2}\in C_{2}$, where $j_{1},j_{2}>2$. Because $\lambda$ is noncrossing, we must have that $j_{1}>j_{2}.$ Therefore, the elements of $C_{1}$ which are not $1$ are greater than all of the elements of $C_{2}$. We see that, where $1\leq j=|C_{1}|\leq|B_{1}|$, $C_{1}$ is the largest j1j-1 elements of $B_{1}$ with $1$ added, and $C_{2}$ is the remaining elements of $B_{1}$ with $2$ added. Therefore, $\lambda\in S(\pi)$.
@@
Assume $\lambda\in S(\pi)$, where $\pi=B_{1}\sqcup\cdots\sqcup B_{m^{\prime}}\in NC^{\prime}(k)$. Since $\lambda$ does not have $2$ in the first block, $\lambda$ is $\pi$ with $B_{1}^{\prime}$ and $B_{1}^{\prime\prime}$ instead of $B_{1}$, where $B_{1}^{\prime}$ is the last j1j-1 elements of $B_{1}$ with $1$ added, and $B_{1}^{\prime\prime}$ is the remaining elements of $B_{1}$ with $2$ added, for some $1\leq j\leq|B_{1}|$. Then, the first and second block of $\lambda$ would be $C_{1}=B_{1}^{\prime}$ and $C_{2}=B_{1}^{\prime\prime}$, respectively. But, in $\pi$, $B_{1}=(B_{1}^{\prime}\cup B_{1}^{\prime\prime})/\{2\}=(C_{1}\cup C_{2})/\{2\}$ and $B_{i}=C_{i+1}$ for $2\leq i\leq m-1$, where $m^{\prime}=m-1$. Thus, $\pi$ is unique. ∎
With ˜2.7.1,
@@
giving Lemma˜2.7 since $|\{1,3,4,\ldots,k+1\}|=|\{1,2,3,\ldots,k\}|=k$. We are done. ∎
@@
The free cumulants appear in (7) and these free cumulants will be used in Section˜3 and Section˜4 to obtain the free cumulants corresponding to the moments $m_{k}$ in Theorem˜1.6 that we mentioned previously. Furthermore, Theorem˜1.8 is based on the asymptotics of the free cumulants.
@@
Let $i$ and $k$ be positive integers. Suppose $\{f_{j,N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ for $1\leq j\leq k-1$ and $\{g_{N}(x_{1},\ldots,x_{N})\}_{N\geq i}$ are sequences of formal series which are symmetric outside of $i$. For $1\leq j\leq k-1$, assume that $f_{j,N}(x_{1},\ldots,x_{N})$ has limiting sequence $f$ outside of $i$. Also, suppose $g_{N}(x_{1},\ldots,x_{N})$ has limiting sequence $g$ outside of $i$. Then,
@@
From Corollary˜2.4, the limit of $\prod_{j=1}^{k-1}\mathcal{Q}_{i}^{N}(f_{j,N})g_{N}(x_{1},\ldots,x_{N})$ as $N\rightarrow\infty$ outside of $i$ exists, see the definition of being symmetric outside of $i$. Thus, the left hand side of (7) is
$1]\lim_{N\rightarrow\infty}\prod_{j=1}^{k-1}\mathcal{Q}_{i}^{N}(f_{j,N})g_{N}(x_{1},\ldots,x_{N}),$
@@
Next, assume that the statement is true for a positve integer $k$. We want to show the statement is true for $k+1$. From Proposition˜2.5, outside of $i$,
@@
Using Corollary˜2.4, let $g^{\prime}$ be the limiting sequence of $\mathcal{Q}_{i}^{N}(F_{i}(f))F_{i}(g)(x_{1},\ldots,x_{N})$ outside of $i$, with
@@
Also, from Lemma˜2.7,
@@
Fix the value of $a$. We use induction on $m$, from $m=a$ to 0. The $m=a$ base case is clear. Suppose the statement holds for $m+1$, $0\leq m<a$. We show the statement holds for $m$ with induction on $b$, where the base case $b=0$ is clear. For the inductive step, assume the statement for $m$, $a$, and $b\geq 0$ holds. We want to show the statement for $m$, $a$, and $b+1$ holds. Note that
@@
Afterwards, the coefficient of $c_{f}^{d_{1},\nu_{1}}c_{g}^{d_{2},\nu_{2}}$ on the left hand side of (9) is, with ˜2.9.1 and ˜2.9.2,
@@ 2.3. Distinct Indices
Also, we look at the constant term after using combinations of $\mathcal{Q}_{i}^{j}$ with distinct indices $i$ in Theorem˜2.13. This is important for the proof of Theorem˜1.6, as will be seen in Section˜3 and Section˜4, and for this, we use the operator in (10). But, it is necessary to consider the free cumulants of $s=\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$ which are symmetric in the $x_{i}$ beforehand.
@@
By Corollary˜2.4, the resulting formal series has variables $x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{j}$ and is symmetric.
@@
Induction on $m$ is used. For the base case $m=1$, $\lim_{N\rightarrow\infty}\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ exists outside of $1$ from Corollary˜2.4 because $s_{1,N}f_{1,N}$ and $s_{1,N}g_{N}$ have limits outside of $1$. For some $i\geq 1$, consider $N\geq i+1$. From Lemma˜2.11, $\mathcal{R}_{N,N}^{\lambda_{1}}(f_{1,N})(g_{N})\in\mathcal{F}_{i}^{N-1}$.
From Proposition˜2.6, $s_{1,N}\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})=\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$. With this, the coefficient $c^{d,\nu}$ of $\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})|_{x_{N}=0}$ as a formal series in $x_{1},\ldots,x_{N-1}$ outside of $i$ will be the coefficient $c^{0,\nu+(d)}$ of $\mathcal{Q}_{N}^{N}(f_{1,N})^{\lambda_{1}}(g_{N})$ outside of $N$, and thus the coefficient $c^{0,\nu+(d)}$ of $\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ outside of $1$. Since the limit of $\mathcal{Q}_{1}^{N}(s_{1,N}f_{1,N})^{\lambda_{1}}(s_{1,N}g_{N})$ outside of $1$ exists, the limit of this coefficient from $N=i+1$ to $\infty$ exists. As this holds for all $d,\nu$, the base case is complete.
@@
By the inductive hypothesis and because $h_{N-m}$ is symmetric by Lemma˜2.11, the limit of $s_{1,N-m}h_{N-m}=h_{N-m}$ from Nm=1N-m=1 to $\infty$ exists outside of $1$. Then,
@@
is the base case $m=1$ with $f_{m+1,N}$ as $f_{1,N-m}$, $h_{N-m}$ as $g_{N-m}$, and NmN-m as $N$, where the conditions are satisfied. Therefore, from $m=1$, the limit from Nm=i+1N-m=i+1 to $\infty$ of the above expression outside of $i$ exists for $i\geq 1$. This limit is from $N=i+m+1$ to $\infty$, which completes the proof. ∎
@@
We prove this with induction on $m$. For the base case $m=1$, use Theorem˜2.9. Suppose the result holds for $m\geq 1$. We want to show it holds for $m+1$.
@@
Then, $s_{1,N-m}\mathcal{Q}_{N-m}^{N-m}\left(f_{m+1,N}\right)^{\lambda_{m+1}}h_{N-m}$ is $\mathcal{Q}_{1}^{N-m}\left(s_{1,N-m}f_{m+1,N}\right)^{\lambda_{m+1}}h_{N-m}$ from Proposition˜2.6. Since the switch does not change the constant,
@@
The limit of this as $N\rightarrow\infty$ is computed. With $f_{N-m}=s_{1,N-m}f_{m+1,N}$ for $N\geq m+1$, $f_{N}$ has $N\rightarrow\infty$ limit $f_{m+1}$ outside of $1$, and from Lemma˜2.12, $h_{N}$ has a limit outside of $1$; let $h$ be the limiting sequence of $\{h_{N}\}_{N\geq 1}$. Then, from Theorem˜2.9 with NmN-m for $N$,
@@
But, as $h_{N}$ is symmetric in $x_{i}$, $1\leq i\leq N$ for $N\geq 1$, $h$ is symmetric. Then, by Proposition˜2.10, $c_{k}(h)=0$, $k\geq 2$, and $c_{1}(h)=c_{h}^{0,0}$. From this,